Sir please tell what is the phase angle between current and voltage when it pass through pure capacitance


1)60


2)90


3)180


4)0


and also tell how to find it through capacitance

Grade:Upto college level

1 Answers

vikas askiitian expert
13 years ago

E = Eosinwt      ..............................1

Eo is the peak voltage

E across capacitor = dq/c

dq/c = Eosinwt

differentiating this equation

(1/c)(dq/dt) = Eowcoswt

 I = Eowc coswt         ...................2

I = Eo/(1/wc) {coswt}

1/wc = Rc (capacitive reactance)

I = Eowc {sin(pi/2 +wt)}         .............3                  [cos@ = sin(pi/2+@) ]

comparing  eq 1 & 3 we see that current leads the voltage by an angle of pi/2

therefore phase difference bw current & voltage across capacitor is 90o .......

 

approve if u like my ans